OpenStudy (anonymous):

x2 + 16x = 0 find the roots of the quadratic equation

OpenStudy (callisto):

x^2+16=0 x(x+16) =0 x=0 or x=-16

OpenStudy (saifoo.khan):

x(x+16) = 0 so zero & -16

OpenStudy (anonymous):

thank you

OpenStudy (saifoo.khan):

you welcome